OpenStudy (anonymous):

what is the solution to the equation? 36x^2=100 A. -9/25 , 9/25 B. -3/5 , 3/5 C. -25/9 , 25/9 D. -5/3 , 5/3

OpenStudy (***[isuru]***):

hi, \[36x ^{2} = 100\]\[x ^{2} = \frac{ 100 }{36 }\]\[x = \pm \sqrt{\frac{ 100 }{ 36 }}\]\[x = \pm \frac{ 10 }{6 } \]\[x= \pm \frac{ 5 }{ 3 }\] which mean.... \[x =\frac{ 5 }{ 3 } \ or \ x=-\frac{ 5 }{ 3 }\] Hoep this will help ya!!!

OpenStudy (anonymous):

thank you :) it helped a lot

OpenStudy (***[isuru]***):

u r welcome!!